OpenStudy (anonymous):

Los Angeles is about 385 miles from San Francisco. How far apart would the cities be on a map with a scale of 1 in. = 25 mi? If necessary, round to the nearest hundredth.

jimthompson5910 (jim_thompson5910):

1 in = 25 mi ... this is given 1/25 in = 1 mi ... divide both sides by 25 385/25 in = 385 mi ... multiply both sides by 385 15.4 in = 385 mi ... Reduce 385 mi = 15.4 in ... flip the equation --------------------------------------------------- Or you can do it this way (distance on map)/(distance in real life) = (distance on map)/(distance in real life) 1/25 = x/385 1*385 = 25x 25x = 385 x = 385/25 x = 15.4

OpenStudy (anonymous):

(385 miles )(1 inch/25 miles) "miles" cancels so you are just dividing 385 by 25 and are left with inches. You can multiply by the factor because it is like multiplying by one.